a sound wave has a wavelength of 0.20m. what is the phase difference between two points on wire which are 0.65m apart


1
Expert's answer
2021-09-03T08:24:56-0400

We can relate the wavelength "\lambda" with the displacement difference between the points "\Delta x" and the phase difference "\Delta \phi" with the following equation:


"\lambda=\cfrac{2\pi \Delta x}{\Delta \phi} \implies \Delta \phi=\cfrac{2\pi \Delta x}{\lambda}"


We substitute and find the phase difference in radians:

"\Delta \phi=\cfrac{2\pi \Delta x}{\lambda}=\cfrac{2\pi (0.65\,m)}{0.20\,m}= 6.5\pi\,rad=1170^{\circ}"


In conclusion, the phase difference "\Delta \phi" is 6.5"\pi" radians or 1170°.


Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.